The perimeter of a rectangle equals the sun of the lengths of its four sides. If the width of a rectangle is 6 inches and its perimeter is 46 inches, find the length of the rectangle.
Answer by ankor@dixie-net.com(22740)

The perimeter of a rectangle equals the sum of the lengths of its four sides.
If the width of a rectangle is 6 inches and its perimeter is 46 inches,
find the length of the rectangle.
:
Let L = the length
:
The perimeter equation:
2L + 2(6) = 46
2L + 12 = 46
2L = 46 - 12
2L = 34
L = 34/2
L = 17 inches, is the length
:
:
Check this
2(17) + 2(6) = 46